# Temporal-Difference Variational Continual Learning

### v2

@@ 2 Related Work
Other derivative works usually blend VCL with architectural and optimization improvements [35, 33, 36, 37, 38, 39] or different posterior modeling assumptions [34, 40, 41, 42]. We specifically highlight UCB [38], which adapts the learning rate according to the uncertainty of the Bayesian model, and UCL [42], which introduces a different implementation for the VCL objective by proposing the notion of node-wise uncertainty. While their contribution are orthogonal to ours, we adopt UCB and UCL as comparison methods to further show that our proposed objective may also be combined with other variational methods and enhance their performance.
@@ 3 Preliminaries
Problem Statement. In the Continual Learning setting, a model learns from a streaming of tasks, which forms a non-stationary data distribution throughout time. More formally, we consider a task distribution $\mathcal{T}$ and represent each task $t\sim\mathcal{T}$ as a set of pairs $\{(\boldsymbol{x}_{t},y_{t})\}^{N_{t}}$, where $N_{t}$ is the dataset size. At every timestep t [^1] We represent each task with the index $t$, which also denotes the timestep in the sequence of tasks., the model receives a batch of data $\mathcal{D}_{t}$ for training. We evaluate the model in held-out test sets, considering all previously observed tasks.
@@ 4.2 From n-Step KL to Temporal-Difference Targets
The proof is in Appendix D. Proposition 4.4 states that the TD($\lambda$)-VCL objective is a sum of discounted TD targets (up to a normalization constant), effectively representing $\lambda$-returns. In parallel, one can show that the n-Step KL Regularization objective, as a particular case, is a simple average of n-Step TD targets. Fundamentally, the key idea behind these objectives is bootstrapping: they build a learning target estimate based on other estimates. Ultimately, the “$\lambda$-target" in Equation 4.2 provides flexibility for bootstrapping by allowing multiple previous estimates to influence the objective.
@@ 5 Experiments and Discussion
Implementation. We use a Gaussian mean-field approximate posterior and assume a Gaussian prior $\mathcal{N}(0,\sigma^{2}\boldsymbol{I})$, and parameterize all distributions as deep networks. For all variational objectives, we compute the KL term analytically and employ Monte Carlo approximations for the expected log-likelihood terms, leveraging the reparametrization trick [45] for computing gradients. We employed likelihood-tempering [33] to prevent variational over-pruning [46]. Lastly, for test-time evaluation, we compute the posterior predictive distribution by marginalizing out the approximated posterior via Monte-Carlo sampling. We provide further detail about architecture and training in Appendix G and our code [^3] https://anonymous.4open.science/r/vcl-nstepkl-5707.
@@
Benchmarks. We evaluate five benchmarks for Continual Learning (CL). First, we introduce three new benchmarks: PermutedMNIST-Hard, SplitMNIST-Hard, and SplitNotMNIST-Hard. These are more challenging versions of traditional CL benchmarks with similar names. They are significantly harder due to two key restrictions. First, the amount of replay memory that any method can use is limited in both dataset size and the number of tasks. As empirically shown in Appendix H, this creates a much more acute scenario of Catastrophic Forgetting. Second, they enforce the adoption of single-head classifiers. As also shown in Appendix H, this requires the model to account for the potential negative transfer learning among tasks, which makes MNIST/NotMNIST-based benchmarks non-trivial for current research. Next, we also evaluate on two other popular CL benchmarks: CIFAR100-10 and TinyImageNet-10. Both benchmarks are very challenging classification problems, particularly in our setting where no pre-trained representations are used. In Appendix J, we detail all benchmark tasks and specific constraints adopted for robust evaluation.
@@ 6 Closing Remarks
The authors thank Panagiotis Tigas for insightful discussions on variational inference. Luckeciano C. Melo acknowledges funding from the Air Force Office of Scientific Research (AFOSR) European Office of Aerospace Research & Development (EOARD) under grant number FA8655-21-1-7017.
Appendix A Derivation of the n-Step KL Regularization Objective
In this Section, we prove Proposition  4.1:
See 4.1
Proof.
Appendix B Derivation of the Temporal-Difference VCL Objective
Lemma B.1.
$\sum_{k=0}^{n-1}\lambda^{k}a=\frac{a(1-\lambda^{n})}{(1-\lambda)}$ (10)
Proof.
Let $s_{n}=\sum_{k=0}^{n}\lambda^{k}a$. Hence,
Now, we prove Proposition  4.2.
See 4.2
Proof.
We can use Lemma B.1 to expand the sum of KL terms:
$=\mathbb{E}_{\pi}[\mathcal{R}(s_{t},a_{t})+\gamma v_{\pi}(s_{t+1})],$
$=\underbrace{\mathbb{E}_{\boldsymbol{\theta}\sim q_{t}(\boldsymbol{\theta})}\Bigg{[}\sum_{i=0}^{n-1}\log{p(\mathcal{D}_{t-i}\mid\boldsymbol{\theta}})]\Bigg{]}}_{\text{Estimated via MC Sampling}}-\underbrace{\mathscr{D}_{KL}(q_{t}(\boldsymbol{\theta})\mid\mid q_{t-n}(\boldsymbol{\theta}))}_{\text{Bootstrapped via past posterior estimations}}+\underbrace{\Bigg{[}\sum_{i=0}^{n-1}\log{Z_{t-i}}\Bigg{]}}_{\text{Constant w.r.t $\boldsymbol{\theta}$}}.$ (18)
Appendix D TD($\lambda$)-VCL is a discounted sum of n-Step TD targets
See 4.4
Proof.
We start by proving the equivalence between Equation 2 and Equation 7:
Now, we show that Equation 4.2 is a discounted sum of n-Step targets:
Appendix E TD-VCL: A spectrum of Continual Learning algorithms
Let us consider the TD-VCL objective in Equation 4.2:
Trivially, if we set $\lambda=0$, assuming $10^{0}=$ 1, it recovers the Vanilla VCL objective, as stated in Equation 3, regardless of the choice of $n$.
[…]
Applying Equations 22 and 23 to TD-VCL objective, we obtain:
which is exactly the N-Step KL Regularization objective.
Appendix F Impact Statement
Appendix G Implementation Details and Reproducibility
Reproducibility. Reported results are averaged across ten different seeds for PermutedMNIST-Hard, SplitMNIST-Hard, and SplitNotMNIST-Hard, and five seeds for CIFAR100-10 and TinyImageNet-10. Error bars represent 95% confidence intervals, while tables show 2-sigma errors up to two decimal places. We execute all experiments using a single GPU RTX 4090. We provide our implementation code for the proposed methods (TD-VCL, TD-UCB, TD-UCL, and n-Step), as well as considered baselines (Batch MLE, Online MLE, VCL, VCL CoreSet, UCB, and UCL) in https://anonymous.4open.science/r/vcl-nstepkl-5707.
Appendix H Hyperparameters
Batch Size 256 256 256 256 256
Max Epochs 100 100 100 100 100
NN Architecture [100, 100] [256, 256] [150, 150, 150, 150] AlexNet AlexNet
Number of Heads 1 1 1 10 10
Learning Rate 1e-3 1e-3 1e-3 1e-3 1e-3
n-Step KL $n$ 5 4 5 5 2
$\beta$ 5e-3 5e-2 5e-2 3e-5 1e-9
TD($\lambda$)-VCL $n$ 8 4 3 10 2
$\lambda$ 0.5 0.8 0.1 0.5 0.1
$\beta$ 1e-3 5e-2 1e-3 1e-5 1e-9
TD($\lambda$)-UCL $n$ 8 4 3 5 2
$\lambda$ 0.5 0.8 0.1 0.8 0.5
$\beta$ 1e-3 5e-2 1e-3 1e-5 1e-7
TD($\lambda$)-UCB $n$ 8 4 3 8 3
$\lambda$ 0.5 0.8 0.1 0.8 0.1
$\beta$ 1e-3 5e-2 1e-3 1e-5 1e-5
Table 5: Hyperparameters for different methods across benchmarks.
VCL $\beta$ 5e-3 5e-3 5e-3 5e-4 1e-5
UCL $\alpha$ 1.0 10.0 0.5 1.0 10.0
$\beta$ 0.001 1.0 0.001 0.001 1.0
$\gamma$ 0.01 1.0 1.0 0.005 0.1
r 0.5 0.5 0.5 0.5 0.5
$\beta_{kl}$ 5e-3 1e-3 1e-5 1e-4 1e-7
UCB $\alpha$ 1.0 1.0 0.1 10.0 100.0
$\beta$ 1e-2 1e-2 5e-2 5e-5 1e-5
Table 6: Hyperparameters for different methods across benchmarks.
Figure 4: A Replay Buffer analysis on the PermutedMNIST. Each curve represents a model re-trained on a buffer composed of “$T$" previous tasks, “$B$" examples of each. Online MLE only considers the current task. Allowing “unlimited" access to previous task data trivializes the CL setting, and a simple MLE baseline is enough to attain strong results. Nevertheless, as we restrict the replay buffer in size and number of tasks, the benchmark becomes substantially more challenging and shows signs of Catastrophic Forgetting.
Appendix J Benchmarks Description
Appendix K Per Task Performance: Additional Results
K.1 SplitMNIST-Hard
K.2 SplitNotMNIST-Hard
K.3 CIFAR100-10
K.4 TinyImageNet-10
Appendix L Hyperparameters Robustness Analysis
L.1 n-Step KL Regularization
L.2 TD($\lambda$)-VCL
Appendix M Full Table Results
CIFAR100-10
t = 2 t = 3 t = 4 t = 5 t = 6 t = 7 t = 8 t = 9 t = 10
TinyImagenet-10
t = 2 t = 3 t = 4 t = 5 t = 6 t = 7 t = 8 t = 9 t = 10
PermutedMNIST-Hard
t = 2 t = 3 t = 4 t = 5 t = 6 t = 7 t = 8 t = 9 t = 10
[…]
SplitMNIST-Hard SplitNotMNIST-Hard
t = 2 t = 3 t = 4 t = 5 t = 2 t = 3 t = 4 t = 5
[…]
CIFAR100-10
t = 2 t = 3 t = 4 t = 5 t = 6 t = 7 t = 8 t = 9 t = 10
[…]
TinyImagenet-10
t = 2 t = 3 t = 4 t = 5 t = 6 t = 7 t = 8 t = 9 t = 10
[…]

### v3

@@ 2 Related Work
Other derivative works usually blend VCL with architectural and optimization improvements [35, 33, 36, 37, 38, 39, 40] or different posterior modeling assumptions [34, 41, 42, 43, 44]. We specifically highlight UCB [38], which adapts the learning rate according to the uncertainty of the Bayesian model, and UCL [43], which introduces a different implementation for the VCL objective by proposing the notion of node-wise uncertainty. While their contribution are orthogonal to ours, we adopt UCB and UCL as comparison methods to further show that our proposed objective may also be combined with other variational methods and enhance their performance.
@@ 3 Preliminaries
Problem Statement. In the Continual Learning setting, a model learns from a streaming of tasks, which forms a non-stationary data distribution throughout time. More formally, we consider a task distribution $\mathcal{T}$ and represent each task $t\sim\mathcal{T}$ as a set of pairs $\{(\boldsymbol{x}_{t},y_{t})\}^{N_{t}}$, where $N_{t}$ is the dataset size. At every timestep t [^1] For notational simplicity, we use the index $t$ to denote both tasks and timesteps. Note that neither the VCL framework nor our proposed methodology requires knowledge of task boundaries, as argued in the Appendix N., the model receives a batch of data $\mathcal{D}_{t}$ for training. We evaluate the model in held-out test sets, considering all previously observed tasks.
@@ 4.2 From n-Step KL to Temporal-Difference Targets
The proof is in Appendix D. Proposition 4.4 states that the TD($\lambda$)-VCL objective is a sum of discounted TD targets (up to a normalization constant), effectively representing $\lambda$-returns. In parallel, one can show that the n-Step KL Regularization objective, as a particular case, is a simple average of n-Step TD targets. Fundamentally, the key idea behind these objectives is bootstrapping: they build a learning target estimate based on other estimates. Ultimately, the $\lambda$-target" in Equation 4.2 provides flexibility for bootstrapping by allowing multiple previous estimates to influence the objective.
@@ 5 Experiments and Discussion
Implementation. We use a Gaussian mean-field approximate posterior and assume a Gaussian prior $\mathcal{N}(0,\sigma^{2}\boldsymbol{I})$, and parameterize all distributions as deep networks. For all variational objectives, we compute the KL term analytically and employ Monte Carlo approximations for the expected log-likelihood terms, leveraging the reparametrization trick [47] for computing gradients. We employed likelihood-tempering [33] to prevent variational over-pruning [48]. Lastly, for test-time evaluation, we compute the posterior predictive distribution by marginalizing out the approximated posterior via Monte-Carlo sampling. We provide further detail about architecture and training in Appendix G and our code [^3] Our code is available at https://github.com/luckeciano/TD-VCL.
@@
Benchmarks. We evaluate five benchmarks for Continual Learning (CL). First, we introduce three new benchmarks: PermutedMNIST-Hard, SplitMNIST-Hard, and SplitNotMNIST-Hard. These are more challenging versions of traditional CL benchmarks with similar names. They are significantly harder due to two key restrictions. First, the amount of replay memory that any method can use is limited in both dataset size and the number of tasks. As empirically shown in Appendix I, this creates a much more acute scenario of Catastrophic Forgetting. Second, they enforce the adoption of single-head classifiers. As also shown in Appendix H, this requires the model to account for the potential negative transfer learning among tasks, which makes MNIST/NotMNIST-based benchmarks non-trivial for current research. Next, we also evaluate on two other popular CL benchmarks: CIFAR100-10 and TinyImageNet-10. Both benchmarks are very challenging classification problems, particularly in our setting where no pre-trained representations are used. In Appendix J, we detail all benchmark tasks and specific constraints adopted for robust evaluation.
@@ 6 Closing Remarks
The authors thank Panagiotis Tigas for insightful discussions on variational inference. Luckeciano C. Melo acknowledges funding from the Air Force Office of Scientific Research (AFOSR) European Office of Aerospace Research & Development (EOARD) under grant number FA8655-21-1-7017. Yarin Gal is supported by a Turing AI Fellowship financed by the UK government’s Office for Artificial Intelligence, through UK Research and Innovation (grant reference EP/V030302/1) and delivered by the Alan Turing Institute.
@@ 80. Guidelines:
Appendix A Derivation of the n-Step KL Regularization Objective
In this Section, we prove Proposition  4.1:
See 4.1
Proof.
Appendix B Derivation of the Temporal-Difference VCL Objective
Lemma B.1.
$\sum_{k=0}^{n-1}\lambda^{k}a=\frac{a(1-\lambda^{n})}{(1-\lambda)}$ (10)
Proof.
Let $s_{n}=\sum_{k=0}^{n}\lambda^{k}a$. Hence,
Now, we prove Proposition  4.2.
See 4.2
Proof.
We can use Lemma B.1 to expand the sum of KL terms:
$=\mathbb{E}_{\pi}[\mathcal{R}(s_{t},a_{t})+\gamma v_{\pi}(s_{t+1})],$
[…]
Appendix D TD($\lambda$)-VCL is a discounted sum of n-Step TD targets
See 4.4
Proof.
We start by proving the equivalence between Equation 2 and Equation 7:
Now, we show that Equation 4.2 is a discounted sum of n-Step targets:
Appendix E TD-VCL: A spectrum of Continual Learning algorithms
Let us consider the TD-VCL objective in Equation 4.2:
Trivially, if we set $\lambda=0$, assuming $0^{0}=1$, it recovers the Vanilla VCL objective, as stated in Equation 3, regardless of the choice of $n$.
Let us develop $(\text{I})$ and $(\text{II})$ separately by applying the L’Hôpital’s rule. First, for $(\text{I})$:
Now, for $(\text{II})$:
Applying Equations 22 and 23 to TD-VCL objective, we obtain:
which is exactly the N-Step KL Regularization objective.
Appendix F Impact Statement
Appendix G Implementation Details and Reproducibility
Reproducibility. Reported results are averaged across ten different seeds for PermutedMNIST-Hard, SplitMNIST-Hard, and SplitNotMNIST-Hard, and five seeds for CIFAR100-10 and TinyImageNet-10. Error bars represent 95% confidence intervals, while tables show 2-sigma errors up to two decimal places. We execute all experiments using a single GPU RTX 4090. We provide our implementation code for the proposed methods (TD-VCL, TD-UCB, TD-UCL, and n-Step), as well as considered baselines (Batch MLE, Online MLE, VCL, VCL CoreSet, UCB, and UCL) in https://github.com/luckeciano/TD-VCL.
Appendix H Hyperparameters
Batch Size 256 256 256 256 256
Max Epochs 100 100 100 100 100
NN Architecture [100, 100] [256, 256] [150, 150, 150, 150] AlexNet AlexNet
Number of Heads 1 1 1 10 10
Learning Rate 1e-3 1e-3 1e-3 1e-3 1e-3
n-Step KL $n$ 5 4 5 5 2
$\beta$ 5e-3 5e-2 5e-2 3e-5 1e-9
TD($\lambda$)-VCL $n$ 8 4 3 10 2
$\lambda$ 0.5 0.8 0.1 0.5 0.1
$\beta$ 1e-3 5e-2 1e-3 1e-5 1e-9
TD($\lambda$)-UCL $n$ 8 4 3 5 2
$\lambda$ 0.5 0.8 0.1 0.8 0.5
$\beta$ 1e-3 5e-2 1e-3 1e-5 1e-7
TD($\lambda$)-UCB $n$ 8 4 3 8 3
$\lambda$ 0.5 0.8 0.1 0.8 0.1
$\beta$ 1e-3 5e-2 1e-3 1e-5 1e-5
Table 5: Hyperparameters for different methods across benchmarks.
VCL $\beta$ 5e-3 5e-3 5e-3 5e-4 1e-5
UCL $\alpha$ 1.0 10.0 0.5 1.0 10.0
$\beta$ 0.001 1.0 0.001 0.001 1.0
$\gamma$ 0.01 1.0 1.0 0.005 0.1
r 0.5 0.5 0.5 0.5 0.5
$\beta_{kl}$ 5e-3 1e-3 1e-5 1e-4 1e-7
UCB $\alpha$ 1.0 1.0 0.1 10.0 100.0
$\beta$ 1e-2 1e-2 5e-2 5e-5 1e-5
Table 6: Hyperparameters for different methods across benchmarks.
Figure 4: A Replay Buffer analysis on the PermutedMNIST. Each curve represents a model re-trained on a buffer composed of “TT" previous tasks, “BB" examples of each. Online MLE only considers the current task. Allowing “unlimited" access to previous task data trivializes the CL setting, and a simple MLE baseline is enough to attain strong results. Nevertheless, as we restrict the replay buffer in size and number of tasks, the benchmark becomes substantially more challenging and shows signs of Catastrophic Forgetting.
Appendix J Benchmarks Description
Appendix K Per Task Performance: Additional Results
K.1 SplitMNIST-Hard
K.2 SplitNotMNIST-Hard
K.3 CIFAR100-10
K.4 TinyImageNet-10
Appendix L Hyperparameters Robustness Analysis
L.1 n-Step KL Regularization
L.2 TD($\lambda$)-VCL
Appendix M Full Table Results
CIFAR100-10
t = 2 t = 3 t = 4 t = 5 t = 6 t = 7 t = 8 t = 9 t = 10
TinyImagenet-10
t = 2 t = 3 t = 4 t = 5 t = 6 t = 7 t = 8 t = 9 t = 10
PermutedMNIST-Hard
t = 2 t = 3 t = 4 t = 5 t = 6 t = 7 t = 8 t = 9 t = 10
UCL 0.97±0.00 0.95±0.01 0.94±0.01 0.92±0.02 0.89±0.02 0.86±0.04 0.83±0.06 0.78±0.09 0.73±0.12
UCB 0.93±0.01 0.93±0.01 0.92±0.01 0.90±0.01 0.89±0.02 0.87±0.02 0.86±0.02 0.85±0.01 0.83±0.02
SplitMNIST-Hard SplitNotMNIST-Hard
t = 2 t = 3 t = 4 t = 5 t = 2 t = 3 t = 4 t = 5
UCL 0.88±0.04 0.68±0.03 0.83±0.03 0.66±0.06 0.71±0.01 0.63±0.04 0.61±0.00 0.52±0.04
UCB 0.85±0.16 0.79±0.12 0.83±0.06 0.75±0.10 0.70±0.08 0.63±0.06 0.61±0.01 0.61±0.05
CIFAR100-10
t = 2 t = 3 t = 4 t = 5 t = 6 t = 7 t = 8 t = 9 t = 10
UCL 0.65±0.03 0.66±0.07 0.64±0.05 0.62±0.04 0.60±0.05 0.60±0.04 0.58±0.02 0.61±0.02 0.62±0.02
UCB 0.65±0.01 0.65±0.02 0.66±0.02 0.66±0.03 0.66±0.03 0.66±0.01 0.65±0.01 0.64±0.01 0.66±0.01
TinyImagenet-10
t = 2 t = 3 t = 4 t = 5 t = 6 t = 7 t = 8 t = 9 t = 10
UCL 0.55±0.02 0.52±0.03 0.52±0.03 0.52±0.02 0.51±0.02 0.50±0.02 0.52±0.01 0.52±0.01 0.50±0.03
UCB 0.52±0.06 0.51±0.04 0.51±0.02 0.50±0.02 0.48±0.04 0.46±0.01 0.45±0.02 0.44±0.03 0.42±0.03
Appendix N Does TD-VCL Assume Knowledge of Task Boundaries?
In this Section, we argue that the TD-VCL objective (and VCL objectives in general) does not require knowledge of task boundaries, and we provide theoretical and empirical evidence for that. The theoretical argument comes from the principle that the Bayesian framework is self-consistent: given a stream of data, the final posterior distribution should be the same regardless of how many Bayesian updates are executed.
Based on that, the key thing is to realize that the number of updates does not need to be equal to the number of tasks. Mathematically, suppose we have a stream of $T$ tasks (represented by $t$). At a particular update $k$, we may consider a Bayesian update that includes data from multiple ($m$) sequential tasks (e.g., from $t_{a}$ to $t_{a+m}$):
$\mathcal{D}_{k}=\bigcup_{j=0}^{m}\mathcal{D}_{j}.$ (24)
Crucially, this does not impose any assumptions on boundaries. Rather, once we decide where to start and end the data stream for the Bayesian update, there could be potentially many tasks included. Under the same assumptions stated in Section 3, we have that:
$p(\mathcal{D}_{k}\mid\theta)=\prod_{j=0}^{m}p(\mathcal{D}_{j}\mid\theta).$ (25)
And, the recursive relationship (Equation 1) also follows:
$p(\theta\mid\mathcal{D}_{1:k})\propto p(\theta\mid\mathcal{D}_{1:k-1})\prod_{j=0}^{m}p(\mathcal{D}_{k}\mid\theta).$ (26)
Finally, following the same variational objective and ELBO derivation, we arrive at
$\mathcal{L}^{k}(\theta)=\mathbb{E}_{\theta\sim q_{k}(\theta)}\left[\sum_{j=0}^{m}\log p(\mathcal{D}_{k}\mid\theta)\right]-\mathcal{D}_{\text{KL}}(q_{k}(\theta)\,\|\,q_{k-1}(\theta)).$ (27)
Therefore, the objective itself does not discriminate or require task boundaries. TD-VCL will estimate the likelihood terms for multiple terms simultaneously, which is something already done while replaying past tasks.
Table 9: StreamingPermutedMNIST-Hard results. We observe no negative impact in the TD-VCL methods in comparison with PermutedMNIST-Hard, suggesting that these methods do not require knowledge of task boundaries.
StreamingPermutedMNIST-Hard
Method $t=10$
Online MLE 0.54±0.09
Batch MLE 0.64±0.09
VCL 0.82±0.05
VCL CoreSet 0.85±0.04
N-Step TD-VCL 0.89±0.02
TD($\lambda$)-VCL 0.89±0.02
Empirical Evidence. We highlight that most benchmarks – including the ones presented in this work – isolate tasks, which makes it convenient to consider one Bayesian update per task. To provide further practical evidence that the method does not require knowledge of boundaries, we present another benchmark called StreamingPermutedMNIST-Hard. This benchmark does not provide any boundary between tasks. From the full data stream of $T$ tasks, we create sequential streams of data where boundaries are placed randomly and provide them to the methods for continual learning. We execute an evaluation after the complete data stream, considering held-out splits composed of all tasks. We report the average accuracy across them, equivalently to the $t=10$ column in Tables 2 and 3. Table 9 shows the empirical results over 10 seeds. We observe no negative impact in the VCL/TD-VCL methods in comparison with PermutedMNIST-Hard. In fact, some methods improved performance, because we are likely replaying the same task into different chunks, alleviating the catastrophic forgetting challenge. The proposed objectives still outperform all other methods.
Appendix O Further Questions
This Appendix presents additional clarification questions aimed at improving the understanding of the proposed method and experiments. These questions were raised during the peer-review process, and we refer to the OpenReview page for the full discussion.
O.1 What is the computational cost associated with TD-VCL?
We analyze the computational cost from three aspects: training, inference, and hyperparameter search.
Training Cost. The training cost arises from the computation of Equations 4 and 4.2, which depend on two components: (I) Monte Carlo estimation of the likelihood term, and (II) the KL regularization term. (I) corresponds to the standard cross-entropy loss for classification, averaged over samples of $\boldsymbol{\theta}$ from the variational distribution. In practice, we approximate this average with a single sample, so the cost is equivalent to standard classification under the MLE objective. (II) is the KL regularization term, which can be computed in closed form. This computation is lightweight since it does not involve data or forward passes through the network. Overall, the training costs of VCL and TD-VCL are nearly identical, as the main bottleneck lies in (I), which is similar in both methods. Additionally, we employ early stopping, which reduces the number of training epochs and thereby lowers computational requirements compared to prior implementations.
Inference Cost. Bayesian inference is approximated by the posterior predictive distribution $p(y^{*}\mid\mathbf{x}^{*},\mathcal{D}_{1:t})=\mathbb{E}_{q_{t}(\theta)}\big[p(y^{*}\mid\theta,\mathbf{x}^{*},\mathcal{D}_{1:t})\big]$, which we estimate via Monte Carlo sampling. The computational cost of this step depends on the number of parameter samples $\theta$. In our experiments, we use a single sample to ensure computational fairness across all baselines. With this choice, inference reduces to a single forward pass through the network, identical to a standard classifier. Using more samples increase the cost proportionally but also improve predictive performance.
Hyperparameter Search Cost. The cost of hyperparameter search can be managed by constraining the computational budget. As described in Appendix H, we restrict all methods (including baselines) to a budget of at most one GPU-day. Our method introduces two additional hyperparameters beyond those of VCL ($n$ and $\lambda$). Results in Appendix L show moderate to good robustness to these choices, suggesting that the search cost could be further reduced if necessary.
O.2 What is the memory cost of TD-VCL?
We analyze memory usage in terms of maintaining the replay buffer and storing the previous posteriors.
Replay Buffer Cost. The buffer has memory complexity $\mathcal{O}(n)$, where $n$ is the $n$-Step hyperparameter. This cost is no greater than that of the core sets used in VCL CoreSet. In our benchmarks, the replay buffer is intentionally limited to at most 200 data points from previously observed tasks, which is negligible compared to the 60,000 data points in the current task. As a result, the replay buffer does not represent a major bottleneck.
Posterior Storage Cost. The storage of posteriors also has memory complexity $\mathcal{O}(n)$. For smaller networks (such as those used in the MNIST and NotMNIST benchmarks), the memory usage is comparable to that of the replay buffer: for example, storing 200 MNIST data points requires about 0.60 MB, while storing the posterior requires about 0.68 MB (assuming float32 precision). Naturally, the memory required for posteriors increases with larger and deeper networks.
O.3 Is maintaining previous posteriors a major bottleneck? Can we optimize this cost?
As noted in the Limitations (Section 6), TD-VCL may increase memory requirements. However, this is not necessarily a major limitation, and we also discuss strategies to reduce memory usage.
Memory Complexity is a function of $n$, which the user controls. The buffer size and number of posteriors are defined by $n$. If memory is a bottleneck, one can control $n$ to satisfy memory constraints. Crucially, $n$ is not always equal to the number of tasks. Our robustness analysis (Appendix L) shows that performance increases monotonically with $n$ up to a level where it may saturate. Therefore, any $n>1$ should be better than vanilla VCL, and, if performance is expected to saturate, one can also set $n$ to be much lower than the number of tasks. The employed hyperparameters (Appendix H) suggest that we can usually assume a value of $n$ that is lower than the number of tasks.
Assume a memory-efficient variational family $\mathcal{Q}$. Since memory may be a challenge for large Bayesian networks, there are alternative architectures, such as last-layer variational methods or bayesian LoRA adapters [59, 12, 52], which approximates the posterior distribution in a fixed number of parameters. These methods drastically reduce the required memory at the cost of expressiveness of the variational family.
Store previous posteriors in cheaper memory alternatives. Since TD-VCL does not use previous posteriors for inference but only for computing the KL regularization, they do not need to occupy GPU memory. In fact, the regularization term could be computed asynchronously on CPU (or even with an external computer) while the GPU is used to generate predictions and estimate the likelihood terms. While implementation is more involved, it allows the use of both CPU/GPU and avoids having previous posteriors in GPU memory, which is usually the bottleneck.
Estimate TD objective with fewer posteriors but covering older timesteps. A corollary of Proposition 4.4 is that we can represent the learning target as any combination of $n$-step TD targets. This means that we may store posteriors at every $m$ steps, instead of every step. In this case, given $T$ tasks, we only store $T/m$ posteriors. Naturally, this leads to a different way of estimating the learning objective, but ensures that it is covering older tasks to prevent catastrophic forgetting.
Lastly, we highlight that there are realistic Continual Learning settings where storing posteriors is not a major bottleneck. For instance, when continually learning on embedded systems with access to a cloud storage. These embedded systems (mobile phones, wearables) usually present limited storage/GPU memory onboard. Some problems require on-the-fly model adaptation and, for privacy-related reasons, the data must be kept on the device for a limited time. Nonetheless, we may upload model snapshots to the cloud without problems (sometimes this upload is required to conduct quality evaluation or audits). A concrete example is an on-device speech recognition model on a smartwatch adapting to a user’s voice. We believe our TD-VCL objective is well-suited for this problem setting.
O.4 When should one use $n$-Step TD-VCL or TD($\lambda$)-VCL?
TD($\lambda$)-VCL is a generalization of $n$-Step TD-VCL. As we presented in Appendix E, TD($\lambda$)-VCL forms a spectrum of CL algorithms, and we recover $n$-Step TD-VCL when $\lambda\rightarrow 1$. Therefore, the "choice" depends on $\lambda$, which controls how much the learning objective should prioritize recent posteriors. If one believes that the most recent posterior retains the knowledge of previous tasks, then a higher $\lambda$ should work better. Otherwise, one should use lower values as past estimates contain information that has not propagated over the recursive updates. In practice, it depends on the continual learning problem and the potential transfer/interference among tasks. The recommendation is to start from TD($\lambda$)-VCL and tune the $\lambda$ hyperparameter.
O.5 What is the impact of Early Stopping in the presented methods?
We perform an ablation study to evaluate the impact of Early Stopping, with results reported in Table 10. We find that removing Early Stopping does not significantly impact the performance of TD-VCL, while VCL shows a slight degradation. The methods most negatively affected are the MLE baselines. This result is expected, since these models lack any form of regularization, and training without Early Stopping leads to overfitting. Notably, experiments without Early Stopping required approximately 5 to 10 times more training time. As discussed in the main paper, Early Stopping substantially reduces computational cost.
[…]
PermutedMNIST-Hard
t = 2 t = 3 t = 4 t = 5 t = 6 t = 7 t = 8 t = 9 t = 10
Online MLE 0.75±0.04 0.58±0.07 0.55±0.07 0.44±0.05 0.42±0.05 0.36±0.04 0.34±0.05 0.31±0.04 0.30±0.04
Batch MLE 0.94±0.01 0.89±0.02 0.82±0.04 0.69±0.03 0.64±0.04 0.55±0.03 0.52±0.03 0.46±0.04 0.45±0.03
VCL 0.96±0.01 0.94±0.01 0.92±0.01 0.89±0.04 0.86±0.04 0.84±0.04 0.80±0.09 0.78±0.11 0.71±0.15
VCL CoreSet 0.96±0.00 0.95±0.01 0.94±0.01 0.92±0.02 0.90±0.02 0.88±0.02 0.85±0.02 0.83±0.03 0.79±0.09
N-Step TD-VCL 0.95±0.00 0.95±0.00 0.94±0.00 0.93±0.00 0.92±0.01 0.92±0.01 0.89±0.01 0.88±0.02 0.85±0.04
TD($\lambda$)-VCL 0.97±0.00 0.97±0.00 0.96±0.01 0.95±0.00 0.94±0.01 0.93±0.01 0.91±0.01 0.91±0.02 0.89±0.02